I need correct answers. If you do not know how to solve it please do not answer, or copy other similar answers. I will rate the correct answer.
Answer #1

Capacitance C of a parallel palte capacitor is given by C = KeoA/d

where A = area = pi r^2,

e0 = constnat = 8.85*10^-12,

d   = distance between the plates,

K = dieelctric constant (=1 for air)

Chareg Q = CV where V = Volatge

so here

if K is introduced

Q = KCV

so if K is introduced Charge increases

and V = Vo/K

voltgae decreases